OpenStudy (anonymous):

I Will Medal And Fan Best Answer! I need explanation AND answer! Fill in the table using this function rule. y = 5x - 4 x y 5 ? 6 ? 7 ? 9 ?

OpenStudy (anonymous):

so pretty much multiplying the numbers by 5 then subtracting by 4 so the first one would be 21

OpenStudy (anonymous):

For each value of x, substitute it into the expression 5x-4 to get the corresponding y-value. For example, for x = 8\[y = 5\left( 8 \right)-4=40-4=36\]